# CRS-Diff: Controllable Generative Remote Sensing Foundation Model

### v2

@@ Abstract
The emergence of diffusion models has revolutionized the field of image generation, providing new methods for creating high-quality, high-resolution images across various applications. However, the potential of these models for generating domain-specific images, particularly remote sensing (RS) images, remains largely untapped. RS images that are notable for their high resolution, extensive coverage, and rich information content, bring new challenges that general diffusion models may not adequately address. This paper proposes CRS-Diff, a pioneering diffusion modeling framework specifically tailored for generating remote sensing imagery, leveraging the inherent advantages of diffusion models while integrating advanced control mechanisms to ensure that the imagery is not only visually clear but also enriched with geographic and temporal information. The model integrates global and local control inputs, enabling precise combinations of generation conditions to refine the generation process. A comprehensive evaluation of CRS-Diff has demonstrated its superior capability to generate RS imagery both in a single condition and multiple conditions compared with previous methods in terms of image quality and diversity. The code is available at https://github.com/Sonettoo/CRS-Diff.
Keywords:
@@
Remote sensing image Diffusion model Controllable generation
@@
Figure 1: Visualisation results of our proposed CRS-Diff. (a) Single Image generation: The first line shows the examples of generated RS images based on only text. (b) Single-conditional image generation: The second line shows the generation results under the control of one additional condition. (c) Multi-conditional image generation: The last line shows the generation results under the control of multiple conditions.
1 Introduction
@@
Diffusion models [10, 36] are a class of probabilistic generative models that turn noise into a representative data sample. Recently, image generation based on diffusion -model [20, 5, 24, 29, 2, 6, 35, 22] has emerged as a hot research topic since the generated images exhibit high quality, e.g., generating realistic images [20, 5, 24, 29], transforming art styles [39, 48], image super-resolution [41, 14, 31], video generation [9], etc. However, most existing diffusion models focus primarily on general image generation, with insufficient exploration in generating specific types of images, such as remote sensing (RS) images.
RS images differ from traditional RGB images in several aspects, particularly in terms of resolution, coverage area, and informational richness. Covering vast geographical areas, these images typically contain a wealth of surface and atmospheric data. These unique characteristics of RS images thus bring a new challenge for RS image generation, i.e., relying solely on textual control is often inadequate for generating RS images of high quality that meet specific requirements. Thus, more controlled conditions for RS image generation are needed to be explored. Currently, there have been several controlled RS image generation models [4, 13]. For example, [4] shows that the pre-trained diffusion models can be conditioned on cartographic data to generate realistic satellite images. DiffusionSat [13] incorporates the associated metadata such as geolocation as conditioning information. However, these models overlook the use of image-related features as control signals. Additionally, these models lack the flexibility of combining any conditions to generate images.
@@
To address these issues, in this paper, we propose CRS-Diff, i.e., a controllable generative remote sensing foundation model. Specifically, a base diffusion model is first trained tailored for the RS domain based on the Stable Diffusion (SD) model[29], which is capable of converting high-precision textual descriptions into RS images as shown in fig. 1. Building upon this, we integrated two additional control signals into the diffusion model to enable the controlled generation of RS images. These two control signals manage the adjustment of the global and local information of the images, which includes seven additional image control conditions (e.g. semantic segmentation masks, road maps, sketch, etc) as shown in Fig.fig. 1(b). Furthermore, incorporating metadata into the timestep embeddings introduces spatial-temporal information, ensuring that the generated RS images are not only visually realistic but also accurately reflect specific geographical and temporal information. Finally, for each control condition, we can manually set whether to choose that condition, thus making our method flexible to combine any conditions to generate images. In summary, the contributions are threefold:
  1. 1.
    We propose a new controllable generative RS foundation model (CRS-Diff).
  2. 2.
    We introduce a new conditional control mechanism to broaden the image generation space and achieve the effective integration of multiple control information.
  3. 3.
@@
    Experimental results have demonstrated the superiority of our CRS-Diff in generating RS imagery that adheres to specific conditions and surpasses previous RS image generation methods in terms of image quality and diversity.
2 Related work
2.1 Text-to-Image Generation
@@
Text-to-image generation has attracted significant attention in the multimodal field, to generate high-definition images corresponding to given textual descriptions. Early research was focused primarily on GANs[27, 28, 46], with text-conditional GANs emerging as the pioneering end-to-end differential architectures from character to pixel level. For instance, Reed et al. introduced the generative adversarial what-where network, capable of generating 128×128 pixel images. Additionally, Reed et al. developed StackGAN, employing stacked generators to produce clear 256 × 256 pixel remote sensing images. However, these models encounter two main challenges: training instability and limited generalization to open-domain scenes. In addition to GAN-based methods, recent studies have transitioned towards autoregressive models or diffusion models, such as DALL-E[24]. Yong et al. [45] utilized modern Hopfield layers[3, 33] for hierarchical prototype learning[16] in text and image embeddings, to extract the most representative prototypes and implement a coarse-to-fine learning strategy, thereby enhancing the realism of generated remote sensing images.
@@
Diffusion Models[10] are generative models that generate new images by gradually transforming an image from a Gaussian noise state to a target image. This process is composed of two main steps: the forward diffusion process and the reverse generation process. Compared to autoregressive models, diffusion models excel in generating more realistic images through a gradual denoising process. Numerous studies have since focused on enhancing the diffusion model, DALLE-2[24] enhances textual guidance capabilities through integration with the CLIP model, while GLIDE[20] explores diverse guidance methodologies. Conversely, Stable Diffusion(SD)[29] augments training data leveraging the capabilities of the diffusion model, thus improving the generation results.
2.2 Controlled diffusion models
@@
The application of Controlled Diffusion Models (CDM)[47, 12, 19, 13, 50, 15] in the field of Text-to-Image (T2I) generation aims to enable users to precisely dictate the content of generated images. Although traditional text-to-image (T2I) diffusion models are capable of generating images from text descriptions, users often experience limited control over the final output. Controlled diffusion models enable users to specify additional generative details, including style, color, and object positioning, through the introduction of enhanced controllable parameters or mechanisms. For example, ControlNet[47], GLIGEN[15] and T2I-Adapter[19] all incorporate additional control networks or control signals on top of the SD weights to enable integrated control of multiple conditions to reduce training costs. Composer trains a large diffusion model from scratch through a new generative paradigm that allows for flexibility in constructing generative conditions, improving controllability, and achieve just the right results.
@@
In particular, ControlNet[47] can be used in conjunction with diffusion models such as DDPM and LDM to augment the controllability and diversity of the generated images. By introducing additional control signals or conditions, such as textual descriptions, image attribute labels, etc., Based on ControlNet, Zhao et al. [50] proposed that Uni-ControlNet is capable of accepting a wide array of additional control signals or combinations of conditions. By merely fine-tuning the adapters while keeping the original SD model unchanged, Uni-ControlNet significantly reduces the training cost, necessitating only two additional adapters for effective outcomes.
In the field of remote sensing, recent work by Espinosa et al. [4] proposed that pre-trained diffusion models can be conditioned on cartographic data to generate realistic satellite images. DiffusionSat[13], has demonstrated the capability to generate high-resolution satellite data utilizing numerical metadata and textual captions. This approach can be extended to address inverse problems, such as multi-spectral input super-resolution. However, DiffusionSat lacks the capability to incorporate input conditions of additional dimensions and enhanced textual control. In contrast, the introduction of an additional control network enables our model to achieve composite control generation under arbitrary conditions and synthesize images with higher precision, attributable to our sophisticated training strategy.
3 Method
@@
We introduce the diffusion model into the field of remote sensing image generation, aiming to enhance generic image generation capabilities for producing more realistic remote sensing images, subsequently introducing an optimized multi-conditional control mechanism that leverages text, image, and other multidimensional information to guide precise image generation and yield high-precision remote sensing images. The construction of the model consists of two steps: initially, text-image pairs are utilized to train the generative diffusion model weights for remote sensing images, building upon the traditional SD framework, and then the combination of multiple conditions (local conditions and global conditions) is implemented through a conditional control network. We will detail the two-stage training process for CRS-Diff, alongside illustrating the implementation of separate decompositions and combinations of training data.
3.1 Single Image Generation
@@
Initially, we employed a diffusion model-based approach for text-to-image generation. This process involves utilizing a frozen variational autoencoder (VAE) encoder and decoder. The purpose is to convert each image $x\in\mathbb{R}^{C\times H\times W}$ into its corresponding latent space variable $z$, thus circumventing the direct learning of the original image’s conditional data distribution given the text conditions $p(x|\tau)$, and instead focuses on learning the feature distribution of the mapped image feature vector $p(z|\tau)$. Text description corresponding to an image is encoded by a CLIP model[23], which subsequently guides the image generation during the denoising process via a cross-attention mechanism[38]. Additionally, a Classifier-free Guidance (CFG) mechanism[11] is introduced:
$\hat{\theta}(z_{t},c)=\omega\theta(z_{t},c)+(1-\omega)\theta(z_{t}),$ (1)
@@
where $z_{t}=a_{t}x_{0}+\sigma_{t}$ and $\omega$ denotes the bootstrap weight. $\hat{\theta}(z_{t},c)$ is the output of the CFG mechanism. It is a weighted sum of the class-conditional output $\theta(z_{t},c)$ and the unconditional output $\theta(z_{t})$.
@@
Textual information serves as the sole guiding factor in this process. Simultaneously, we employ the pretrained CLIP ViT-L/14@336px model[29], fine-tuned on the RSICD remote sensing image dataset, to amplify the effect of textual guidance. Under the original CFG framework, conditional noise prediction $\theta(x_{t},c)$ is solely dependent on the model’s processing of a given condition (c) (e.g., text description). With the introduction of the CLIP bootstrap, this conditional prediction becomes further influenced by the similarity loss between the image and text as computed by the CLIP model. This implies that the calculation of $\theta(x_{t},c)$ incorporates considerations for better aligning the resulting image with the textual description.
@@
Our text-to-image generation methodology initially encodes an image $x\in\mathbb{R}^{C\times H\times W}$ utilizing a static VAE, which is then converted into a latent representation $z=E(x)\in\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$. Subsequently, Gaussian noise is added to the latent image features to produce a noisy latent representation $z_{t}=\alpha_{t}z+\sigma_{t}\varepsilon$, where $\varepsilon$ represents the Gaussian noise component, and $\alpha_{t}$ and $\sigma_{t}$ are coefficients modulating the noise intensity. The text caption $\tau$ is encoded by a CLIP model $T_{\theta}(\tau)$, generating the text embedding $\tau^{\prime}$. This is accomplished through a denoising model $\epsilon_{\theta}(z_{t};\tau^{\prime},c)$. Finally, the denoised potential representation is up-sampled to the original image resolution using the VAE decoder, completing the image generation process. In this process, we use a sampling algorithm, Denoising Diffusion Implicit Models (DDIM), which is used to speed up the sampling process of diffusion models.
@@
Additionally, the weights of the encoder $E$, decoder $D$, CLIP text encoder $T_{\theta}$, and denoising model $\epsilon_{\theta}$ were inherited from the SD 1.5 version. During the training process, only the denoising model $\epsilon_{\theta}$ is updated to accelerate training convergence and exploit the rich semantic information in the SD model.
3.2 Image Decoupling

### v3

@@ Abstract
The emergence of generative models has revolutionized the field of remote sensing (RS) image generation. Despite generating high-quality images, existing methods are limited in relying mainly on text control conditions and thus don’t always generate images accurately and stablely. In this paper, we propose CRS-Diff, a new RS generative foundation framework specifically tailored for RS image generation, leveraging the inherent advantages of diffusion models while integrating more advanced control mechanisms. Specifically, CRS-Diff can simultaneously support text-condition, metadata-condition, and image-condition control inputs, thus enabling more precise control to refine the generation process. To effectively integrate multiple condition control information, we introduce a new conditional control mechanism to achieve multi-scale feature fusion, thus enhancing the guiding effect of control conditions. To our knowledge, CRS-Diff is the first multiple-condition controllable generative RS foundation model. Experimental results in single-condition and multiple-condition cases have demonstrated the superior ability of our CRS-Diff to generate RS images both quantitatively and qualitatively compared with previous methods. Additionally, our CRS-Diff can serve as a data engine that generates high-quality training data for downstream tasks, e.g., road extraction. The code is available at https://github.com/Sonettoo/CRS-Diff.
Index Terms:
@@
Remote sensing image, Diffusion model, Controllable generation, deep learning
Figure 1: (a) Comparison between natural image and remote sensing (RS) image. The image content is the Capital Museum of China, sourced from Google Maps and Google Street View, respectively. As can be seen, RS imagery differs significantly from traditional RGB imagery in resolution, coverage area, and information richness. (b) Comparison of the generation results between the two control modes. The upper image is the generation result guided solely by text, while the lower image is the result guided by both text and sketch. As can be seen, single text control condition fails to generate accurate image content while ”text + sketch” conditions can succeed.
@@
Figure 2: Visualisation results of our proposed CRS-Diff. (a) Singe text condition generation: the RS images are generated based only on text. (b) Single image condition generation: the RS images are generated based on the image condition. (c) Multi-condition image generation: the RS images are generated under the control of multiple conditions.
I Introduction
@@
Diffusion models [1, 2] are a class of probabilistic generative models that turn noise into a representative data sample. Recently, image generation based on diffusion model [3, 4, 5, 6, 7, 8, 9, 10] has emerged as a hot research topic since the generated images exhibit high quality, e.g., generating realistic images [3, 4, 5, 6], transforming art styles [11, 12], image super-resolution [13, 14, 15], video generation [16], etc. However, most existing diffusion models focus primarily on general image generation, with insufficient exploration in generating specific types of images, such as remote sensing (RS) images.
TABLE I: Comparison of existing controllable generative RS models.
Method Fine-Tuning Text Metadata Image Composable
Control Control Control Control
RSDiff [17]
SatDM [18]
RSFSG [19]
MetaEarth [20]
DiffusionSat [21]
CRS-Diff (Ours)
As shown in Fig. 1 (a), the RS imagery differs significantly from the traditional RGB imagery in several ways, particularly in terms of resolution, coverage area, and information richness. The resolution of remotely sensed images is often very high to capture subtle surface features, unlike the standard resolution of traditional RGB images. In addition, RS images cover wide geographic areas and various environments, such as urban, rural, forest, and marine environments, providing extensive geographic information. In contrast, traditional RGB images usually capture only local areas, offering far less information richness than RS images. Therefore, the high resolution of RS images requires that generative models possess higher accuracy and detail capture capabilities, handle larger scale data, and maintain geographic information consistency during generation. Moreover, the rich information in RS images requires generative models to integrate and represent complex multidimensional data, as relying solely on simple textual control is often insufficient. As shown in the upper part of Fig. 1 (b), textual descriptions, while providing some contextual information, are often inadequate for handling complex geographic and atmospheric data, making it difficult to accurately control the quality and content of the generated images. When generating scene details and regular buildings in the images, although the text-guided images exhibit largely similar features, they often contain distorted line segments and incomprehensible details that contradict the physical world. Therefore, we believe that more fine-grained conditional control is necessary to generate RS images. As shown in the lower part of Fig. 1 (b), incorporating additional conditions (e.g., image sketch) into the image generation process enables the creation of more realistic images. Establishing this correspondence between conditions and images can expand the application scenarios of the generated model. Thus, additional control conditions need to be explored for RS image generation.
Currently, the research in RS image generation mainly includes GAN-based [22, 23, 24, 25, 26] and diffusion model-based approaches [19, 17, 27, 21]. For example, Reed et al. developed StackGAN [24], which employs stacked generators to produce clear RS images with the size of 256 $\times$ 256. However, GAN-based methods are unstable in the training process. In contrast, diffusion models exhibit superior generative ability and a relatively stable training process. As shown in Tab. I, there have been several controlled RS image generation models [28, 19, 21, 20, 18, 17, 29]. For example, RSDiff [17] proposes a novel cascade architecture for RS text-to-image generation using diffusion model. SatDM [18] emphasizes the crucial role of semantic layouts in generating RS images and can produce RS images guided by semantic masks. Yuan et al. [19] notably generate high-quality RS images guided by semantic masks and introduce a lightweight diffusion model, obtained through a customized distillation process, to achieve fast convergence, addressing the inherent issue of prolonged training times in diffusion models. Recently, Yu et al. [20] proposed a guided self-cascading generation framework employing a novel noise sampling strategy, capable of generating images with diverse geographic resolutions across any region for downstream tasks. DiffusionSat [21] incorporates the associated metadata such as geolocation as conditioning information to generate the RS image. However, these models lack control over the image detail level, still using text as the primary control condition and neglecting to incorporate image-related features as control signals. A single text-guided generated image can easily suffer from partial distortion (as shown in Fig. 1 (b)), making it difficult to adapt to the high information density of RS images, rendering the generated images of limited use to downstream tasks.
@@
To address these issues, in this paper, we propose CRS-Diff, i.e. a controllable generative remote sensing foundation model. Specifically, a base diffusion model is first trained tailored for the RS domain based on the Stable Diffusion (SD) model [6], which is capable of converting high-precision textual descriptions into RS images as shown in Fig. 2 (a). Based on this, we integrated ControlNet [30] to include two additional control signals in the diffusion model for the controlled generation of RS images. These two control signals adjust global and local condition information of the image, including six additional image control conditions (semantic segmentation mask, roadmap, sketch, etc.) and textual conditions (prompt, content image, and metadata encoding) as shown in Fig. 2 (b). The optional combination of multiple conditions is controlled to ensure that the resulting RS images are visually realistic and accurately reflect specific geographic and temporal information. For the text condition, we concatenate directly with the original text encoding through an additional encoding step, leveraging the model’s natural control mechanism. For the image condition, we explore multiscale feature fusion to coordinate different control conditions and efficiently implement the bootstrapping of generative process noise maps, making our method flexible enough to combine any conditions for image generation, as shown in Fig. 2 (c).
In summary, the contributions of our work are threefold:
  • We propose a new controllable generative RS foundation model (CRS-Diff), which is a foundational framework specifically designed for RS image generation. Different from previous RS generative models, our CRS-Diff can simultaneously support more types of controllable conditions, i.e., text, metadata and image.
  • To effectively integrate multiple control information, we introduce a new conditional control mechanism to achieve multi-scale feature fusion to enhance the guiding effect of control conditions, thus broadening the image generation space. As far as we know, our CRS-Diff is the first multiple-condition controllable generative RS foundation model, which is capable of generating high-quality RS images that meet specific requirements under the guidance of composite conditions.
@@
  • Experimental results have demonstrated the superiority of our proposed CRS-Diff in generating RS imagery that adheres to specific conditions and surpasses previous RS image generation methods both quantitatively and qualitatively. Additionally, our CRS-Diff can serve as a data engine that generates high-quality training data for downstream tasks, e.g., road extraction.
The rest of this paper is organized as follows. Section II provides a brief overview of related work. Section III details the technical aspects of the CRS-Diff implementation. Section IV describes the experimental design, presents the experimental results, and offers specific analyses. Finally, Section V presents the conclusions of the paper.
II Related work
II-A Text-to-Image Generation
@@
Text-to-image generation that generates high-definition images corresponding to given textual descriptions has attracted significant attention in the multimodal field. Early research is primarily focused on GANs [31, 32, 33], with text-conditional GANs emerging as pioneering end-to-end differential architectures from character to pixel level. For example, Reed et al. [32] introduced the generative adversarial network, capable of generating 128$\times$ 128-pixel images. Additionally, Zhang et al. [33] developed StackGAN, employing stacked generators to produce clear 256$\times$256 pixel RS images. However, these models face two main challenges: training instability and limited generalization to open-domain scenes.
In addition to GAN-based methods, recent studies have shifted toward autoregressive models for text-to-image generation, using web-scale image-text pairs, such as DALL-E [5]. These models demonstrate robust generative capabilities, particularly in zero-shot settings for open-domain scenes, starkly contrasting the small-scale data focus of GAN-based approaches. OpenAI’s DALL-E, leveraging large transformer models and extensive training data, effectively maps language concepts to the pixel level, generating high-quality 256$\times$256 images. Furthermore, Yong et al. employed modern Hopfield layers [34, 35] for hierarchical prototype learning [36] in text and image embeddings, aiming to extract the most representative prototypes and implement a coarse-to-fine learning strategy. These prototypes are then used to encapsulate more complex semantics in text-to-image tasks, enhancing the realism of generated RS images.
@@
Diffusion Models [1] are generative models that generate new images by gradually transforming an image from a Gaussian noise state to a target image. This process contains two main steps: the forward diffusion process and the reverse generation process. Compared to autoregressive models, diffusion models excel in generating more realistic images through a gradual denoising process. Numerous studies have since focused on enhancing the diffusion model, DALLE-2 [5] enhances textual guidance capabilities through integration with the CLIP model, while GLIDE [3] explores diverse guidance methodologies. Conversely, Stable Diffusion (SD) [6] augments the training data by leveraging the capabilities of the diffusion model, thus improving the generation results.
II-B Controlled diffusion models
@@
Controlled Diffusion Models (CDM) [30, 37, 38, 21, 39, 40] for text-to-image (T2I) generation aims to enable users to precisely dictate the content of generated images. Although traditional T2I models can generate images from text descriptions, users often experience limited control over the final output. Controlled diffusion models enable users to specify additional generative details, including style, color, and object positioning, through the introduction of enhanced controllable parameters or mechanisms. For example, ControlNet [30], GLIGEN [40] and T2I-Adapter [38] incorporate additional control networks or control signals on top of the weights of SD to enable integrated control of multiple conditions to reduce training costs. The composer trains a large diffusion model from scratch through a new generative paradigm that allows flexibility in the construction of generative conditions, improves controllability, and achieves better results.
@@
In particular, ControlNet [30] can be used in conjunction with diffusion models such as Denoising Diffusion Probabilistic Models (DDPM) to augment the controllability and diversity of the generated images. By introducing additional control signals or conditions, such as textual descriptions, image attribute labels, etc., Zhao et al. [39] proposed Uni-ControlNet, which supports various additional control signals or combinations of conditions. By fine-tuning the adapters while keeping the original SD model unchanged, Uni-ControlNet significantly reduces the training cost, necessitating only two additional adapters for effective outcomes.
In the remote sensing field, many RS generative models have been proposed. For example, Espinosa et al. [28] proposed a pre-trained diffusion model that is conditioned on cartographic data to generate realistic satellite images. RSDiff [17] introduces a new architecture consisting of two cascading diffusion models for RS text-to-image generation. Yuan et al. [19] introduced a lightweight diffusion model obtained through a customized distillation process, which enhances the quality of image generation via a multi-frequency extraction module and achieves fast convergence by resizing the image at different stages of the diffusion process. SatSynth [29] can simultaneously generate images and corresponding masks for satellite image segmentation, which can then be applied to data augmentation. DiffusionSat [21], has demonstrated the capability to generate high-resolution satellite data utilizing numerical metadata and textual captions. In contrast, we employed a sophisticated training strategy to design and introduce an additional control network, enabling our model to achieve composite control generation under various conditions.
Figure 3: The overall architecture of our proposed CRS-Diff model. CRS-Diff is mainly based on Stable Diffusion (SD), that diffusion process is performed in latent space. The training of CRS-Diff contains two stages. In the first stage of the training process, we train the backbone U-Net network of SD on text-image pairs. The diffusion network obtained from this training is frozen (blue area) during the second training phase, and the encoder and intermediate blocks are copied into ControlNet to adapt to conditional inputs. In the second stage of training, we stack the conditional images as inputs and extract conditional features using a feature extractor. These features are gradually injected into the encoder of ControlNet (orange area) through a Feature Fusion (FF) module. Here, we use a convolutional network to reshape the obtained feature vectors to the current noise dimension and then integrate them with the noise output of the current block of the ControlNet encoder through Attention Feature Fusion (AFF), achieving multi-scale conditional injection.
III Method
@@
We introduce the diffusion model into the field of remote sensing image generation, aiming to enhance generic image generation capabilities for producing more realistic remote sensing images, subsequently introducing an optimized multi-conditional control mechanism that leverages text, image, and other multidimensional information to guide precise image generation and yield high-quality RS images. The construction of the model consists of two steps: initially, text-image pairs are utilized to train the generative diffusion model weights for RS images, building upon the traditional SD framework, and then the combination of multiple conditions (image conditions and text conditions) is implemented through a conditional control network. We will detail the two-stage training process for CRS-Diff, alongside illustrating the implementation of separate decompositions and combinations of training data.
III-A Text-to-Image generation
@@
Initially, we employed the Stable Diffusion (SD) [6] for text-to-image generation. This process involves utilizing a frozen variational autoencoder (VAE) encoder and decoder. The purpose is to convert each image $x\in\mathbb{R}^{C\times H\times W}$ into its corresponding latent space variable $z$, thus circumventing the direct learning of the original image’s conditional data distribution given the text conditions $p(x|\tau)$, and instead focuses on learning the feature distribution of the mapped image feature vector $p(z|\tau)$. Text description corresponding to an image is encoded by a CLIP model [41], which subsequently guides the image generation during the denoising process via a cross-attention mechanism [42]. Thus, the training process of the diffusion model involves updating the latent space related to the U-Net. By predicting the noise added in the forward process and removing it in the reverse process, the model can learn the data distribution in the latent space. The training objective in this process is defined as follows:
$\min_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{z,\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,c)\|_{2}^{2}\right],$ (1)
where $\theta$ is the parameters of the model being optimized, $\epsilon$ is the noise added in the forward process, $\epsilon_{\theta}(z_{t},t,c)$ denotes the prediction of the noise given the noisy data $z_{t}$, time $t$, and condition $c$.
Additionally, a Classifier-free Guidance (CFG) [43] mechanism is introduced:
$\hat{\epsilon}_{\theta}(z_{t},c)=\omega\cdot{\epsilon}_{\theta}(z_{t},c)+(1-\omega)\cdot{\epsilon}_{\theta}(z_{t}),$ (2)
@@
where $z_{t}=a_{t}z_{0}+\sigma_{t}$ and $\omega$ denote the bootstrap weight. $\hat{{\epsilon}_{\theta}}(z_{t},c)$ is the output of the CFG mechanism. It is a weighted sum of the class-conditional output ${\epsilon}_{\theta}(z_{t},c)$ and the unconditional output ${\epsilon}_{\theta}(z_{t})$.
Figure 4: The raw text was encoded using a CLIP text encoder fine-tuned on RS images. The content image is initially encoded using the CLIP image encoder, and the resulting encoding is then converted into four additional text tokens by a Feed Forward Network(FFN). The metadata is first mapped into fixed intervals and then converted into the same number of tokens by an embedding layer. Finally, these processed encodings are concatenated, replacing the original text-encoded input.
@@
Textual information serves as the sole guiding factor in this process. Simultaneously, we employ the pre-trained CLIP ViT-L/14@336px model [6], fine-tuned on the RSICD RS image dataset, to amplify the effect of textual guidance. Under the original CFG framework, conditional noise prediction ${\epsilon}_{\theta}(z_{t},c)$ is solely dependent on the model’s processing of a given condition (e.g., text description). With the introduction of the CLIP bootstrap, this conditional prediction becomes further influenced by the similarity loss between the image and text as computed by the CLIP model. This implies that the calculation of ${\epsilon}_{\theta}(z_{t},c)$ incorporates considerations for better aligning the resulting image with the textual description.
@@
Our text-to-image generation methodology initially encodes an image $x\in\mathbb{R}^{C\times H\times W}$ utilizing a static VAE, which is then converted into a latent representation $z=E(x)\in\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$. Subsequently, Gaussian noise is added to the latent image features to produce a noisy latent representation $z_{t}=\alpha_{t}z+\sigma_{t}\varepsilon$, where $\varepsilon$ represents the Gaussian noise component, and $\alpha_{t}$ and $\sigma_{t}$ are coefficients modulating the noise intensity. The text caption $\tau$ is encoded by a CLIP model $T_{\theta}(\tau)$, generating the text embedding $\tau^{\prime}$. This is accomplished through a denoising model $\epsilon_{\theta}(z_{t};\tau^{\prime},c)$. Finally, the denoised potential representation is up-sampled to the original image resolution using the VAE decoder, completing the image generation process. In this process, we use the Denoising Diffusion Implicit Models (DDIM) algorithm to sample, which can speed up the sampling process of diffusion models.
@@
Additionally, the weights of the encoder $E$, decoder $D$, CLIP text encoder $T_{\theta}$, and denoising model $\epsilon_{\theta}$ are inherited from the SD 1.5 version. During the training process, only the denoising model $\epsilon_{\theta}$ is updated to accelerate training convergence and exploit the rich semantic information in the SD model.
III-B Image Decoupling